What is the average atomic mass for copper, if it is made of copper-62 which is 55% abundant, copper-65 which is 40% abundant and copper-66 which is 5% abundant
1
Expert's answer
2020-11-20T09:56:07-0500

The average atomic mass of an element is the sum of the masses of its isotopes, each multiplied by its natural abundance.

"Avg = (62 \times 0.55) + (65 \times 0.40) + (66 \times 0.05) = 34.1 + 26.0 + 3.3 = 63.4 \; amu"